# XAI4LLM. Let Machine Learning Models and LLMs Collaborate for Enhanced In-Context Learning in Healthcare

### v1

@@ 2.1. Generative models and LLMs
    Pretrained Generative Models. These models utilize the vast knowledge acquired from being trained on a broad spectrum of data, textual, visual (and possibly tabular). By fine-tuning these models on disease-specific or demographic-specific datasets, their capacity to accurately interpret complex medical scenarios and patient histories is enhanced, thus improving their ability to offer precise diagnostic insights or personalized treatment plans. They can be used in the following scenarios (Ge et al., 2023; Moor et al., 2023; Spolaôr et al., 2024; Henriksson et al., 2023; Shentu and Al Moubayed, 2024):
@@
      Embedding Generation for Downstream Training. This technique is particularly useful for interpreting complex medical data such as electronic health records or imaging data, where embeddings can capture intricate patterns and relationships (Yang et al., 2023).
@@ 2.3. LLMs for Automated Healthcare Diagnosis
The integration of LLMs into healthcare has been an area of significant research interest, aiming to enhance diagnostic accuracy and decision-making processes. The work at hand contributes to advancing the application of LLMs in medical diagnostics, each adopting unique approaches to harness the capabilities of LLMs in addressing complex healthcare challenges. A few recent research works are reviewed in the following:
@@
The integration of LLMs into healthcare diagnostics is an area of significant research interest, aiming to enhance both diagnostic accuracy and decision-making processes. Our paper contributes to advancing the application of LLMs in medical diagnostics by proposing a systematic approach that involves structured prompts integrating medical knowledge into the decision-making process, thus enhancing the capabilities of LLMs in analyzing patient data. In a previous version of this work (Nazary et al., 2023), we leveraged ChatGPT alongside few-shot and zero-shot ICL to enhance decision-making in healthcare. This approach focused on extracting and incorporating domain knowledge from trained ML models into prompt templates for healthcare applications, aiming to harness the adaptability of ChatGPT for specific challenges in healthcare decision support. The current work at hand can be seen as an extension of our previous work but does this in a much more systematic way, and explores various unlooked or overlooked areas, such as optimized integration of domain-specific knowledge into prompts, and the design of systematic structured prompt templates, optimizing communication styles between LLMs and users for enhanced diagnostic accuracy. It also addresses the challenge of processing and interpreting numerical data within medical contexts, and evaluating the inherent risks such as false negatives and biases in LLM-based clinical decision-making systems.
@@ 5. Experimental Configuration
API Model of OpenAI. For assessing the risk of heart disease from patient health records, we integrated OpenAI GPT-3.5-turbo model. This model facilitates interaction through a conversational interface, processing a sequence of messages and user inputs as prompts to produce relevant responses. The model temperature was set to 0, ensuring deterministic output behavior. Utilization of the OpenAI model included batching requests containing patient data features, with each batch processed to yield a prediction. The implementation leveraged the OpenAI API, the tenacity library for retrying upon failure, along with standard data processing libraries. For replication purposes, the source code has been made accessible. [^5] Source code available at https://l1nk.dev/EbrXw

### v2

@@ 2.1. Generative models and LLMs
    Pretrained Generative Models. These models utilize the vast knowledge acquired from being trained on a broad spectrum of data, textual, visual (and possibly tabular). By fine-tuning these models on disease-specific or demographic-specific datasets, their capacity to accurately interpret complex medical scenarios and patient histories is enhanced, thus improving their ability to offer precise diagnostic insights or personalized treatment plans. They can be used in the following scenarios (Ge et al., 2023; Moor et al., 2023; Spolaôr et al., 2024; Henriksson et al., 2023; Shentu and Al Moubayed, 2024; He et al., 2023):
@@
      Embedding Generation for Downstream Training. This technique is particularly useful for interpreting complex medical data such as electronic health records or imaging data, where embeddings can capture intricate patterns and relationships (Yang et al., 2023; Li et al., 2023).
@@ 2.3. LLMs for Automated Healthcare Diagnosis
The integration of LLMs into healthcare has been an area of significant research interest, aiming to enhance diagnostic accuracy and decision-making processes (Sayin et al., 2024; Kusa et al., 2023; Galitsky, 2024; Goyal et al., 2024). The work at hand contributes to advancing the application of LLMs in medical diagnostics, each adopting unique approaches to harness the capabilities of LLMs in addressing complex healthcare challenges. A few recent research works are reviewed in the following:
@@
This work builds upon and extends considerably a work we presented at the XI-ML@ECAI workshop (Nazary et al., 2023). In this work, we prototyped the use of zero-shot and few-shot learning for in-context learning (ICL) in the medical domain. We integrated domain knowledge using feature-level explanations from various ML models and demonstrated that this approach, combined with few-shot examples, allows the LLM to achieve quality comparable, though slightly lower, to that of traditional ML models. The promising results motivated further exploration. Building on this foundation, our current work extends the research in several key areas. We developed a comprehensive framework for ICL in the medical domain, formalizing a multi-layer prompt design for clinical decision-making. We explored chain-of-thought versus direct request approaches, and most importantly introduced a new study that highlights the importance of information flow from the prompt construction layer to the LLM. Our findings indicate that single-turn approaches reach peak accuracy with fewer shots but are sensitive to longer prompts, whereas conversational approaches are less sensitive and perform better with larger shots. Additionally, we released code for automatically generating patient prompts based on demographics, improved domain knowledge integration using quartiles, and expanded our evaluation to include risks such as false negatives and gender bias, discovering nuanced aspects like LLMs being more biased than ML models though few-shot learning helps to mitigate this bias.
@@ 5. Experimental Configuration
API Model of OpenAI. For assessing the risk of heart disease from patient health records, we integrated OpenAI GPT-3.5-turbo model. This model facilitates interaction through a conversational interface, processing a sequence of messages and user inputs as prompts to produce relevant responses. The model temperature was set to 0, ensuring deterministic output behavior. Utilization of the OpenAI model included batching requests containing patient data features, with each batch processed to yield a prediction. The implementation leveraged the OpenAI API, the tenacity library for retrying upon failure, along with standard data processing libraries. For replication purposes, the source code has been made accessible. [^5] Source code available at https://shorturl.at/ayNRV (core code), and https://shorturl.at/iowE9 (evaluation code).